OpenStudy (anonymous):

Suppose y varies directly with x. IF y=6 when x=2, find x when y=15

OpenStudy (anonymous):

if y varies directly with x, then that means y = k*x where k is a constant. so you have 6 = k*2, you can solve for k. Then for when y = 15 you will have 15 = k*x and you should have just solved for k , so you can then solve for x.